# Large Language Models for Software Engineering: A Systematic Literature Review

### v5

@@ 6.7. How are LLMs used in software management?
RQ4 - Summary (1) We categorized SE tasks into six activities: requirements engineering, software design, software development, software quality assurance, software maintenance, and software management. Subsequently, we summarized the specific applications of LLMs in these SE activities. (2) We identified a total of 55 SE tasks and found that LLMs are most widely used in software development, with 136 papers mentioning 21 SE tasks. The least applied area, software management, was mentioned in only one study. (3) Code generation and program repair are the most prevalent tasks for employing LLMs in software development and maintenance activities. We analyze the top-performing LLMs repeatedly validated in these tasks and summarize novel findings.
@@ 7. Threats to Validity
Paper search omission. One key limitation is the possibility of omitting relevant papers during the search process. When gathering papers related to LLM4SE tasks from various publishers, it is possible to miss some papers due to incomplete summarization of keywords for software engineering tasks or LLMs. To address this concern, we adopted a comprehensive approach, combining manual search, automated search, and snowballing techniques, to minimize the risk of missing relevant papers. For the manual search, we diligently searched for LLM papers related to SE tasks in six top-tier SE venues and extracted authoritative and comprehensive SE tasks and LLM keywords from these sources. With these numbered keyword search strings in place, we conducted automated searches on seven widely used publisher platforms. Additionally, to further augment our search results, we employed both forward and backward snowballing.
Study selection bias. Another limitation is the potential study selection bias. We established inclusion and exclusion criteria to perform the initial selection of papers, followed by manual verification based on quality assessment criteria (QAC). This process involves a combination of automated and manual procedures. The automated selection process may result in mislabeling of papers due to incomplete or ambiguous information in their corresponding BibTeX records. To mitigate this issue, any papers that cannot be confidently excluded are temporarily retained for manual verification. However, the manual verification stage could be influenced by the subjective judgment biases of the researchers, affecting the accuracy of the quality assessment of papers. To address these concerns, we invited two experienced reviewers in the fields of SE and LLM research to conduct a secondary review of the study selection results. This step aims to enhance the accuracy of our paper selection and minimize the likelihood of omission or misclassification. By implementing these measures, we strive to ensure that the selected papers are accurate and comprehensive, minimizing the impact of study selection bias and enhancing the reliability of our systematic literature review. We additionally provide a replication package [^5] https://docs.google.com/spreadsheets/d/1iomMvoDL2znNDQ_J4aGnqb3BhZpEMlfz for others to view.
@@ 8.1. Challenges
Model size and deployment. The size of LLMs has seen a marked increase over time, moving from GPT-1’s 117M parameters to GPT-2’s 1.5B, and further to GPT-3’s 175B parameters (Yang et al., 2023a). The billions and even trillions (Moss, 2021) of parameters pose significant storage, memory, and computational challenges, which can hinder LLMs in resource-limited and real-time scenarios, especially when developers lack access to powerful GPUs or TPUs. CodeBERT (Feng et al., 2020), a pre-trained model proposed in 2019, has a total of 125M parameters, resulting in a large model size of 476 MB. Recently proposed models like Codex (Chen et al., 2021b) and CodeGen (Nijkamp et al., 2022a), have over 100 billion parameters and over 100 GB in size. The large sizes also require more computational resources. As pointed out by Hugging Face team (Bekman, 2022), training a 176B model (i.e., BLOOM (Scao et al., 2022)) on 1.5 TB datasets consumes an estimated 1,082,880 GPU hours. Similarly, the training of the GPT-NeoX-20B model (Black et al., 2022) on the Pile dataset (Gao et al., 2020), encompassing over 825 GiB of raw text data, requires the deployment of eight NVIDIA A100-SXM4-40GB GPUs. Each of these GPUs comes with a price tag of over 6,000 dollars (Amazon, 2023b), and the training extends to 1,830 hours or approximately 76 days. Moreover, even training a relatively smaller model like the PolyCoder (2.7B) (Xu et al., 2022), employing eight NVIDIA RTX 8000 GPUs on a single machine, demands a commitment of around 6 weeks. These examples illustrate the significant computational costs associated with training LLMs. These also have significant energy costs with predictions of massively increased energy usage by LLM-based platforms (Rillig et al., 2023). Fortunately, there are preliminary studies on reducing code models’ size and improving their efficiency. Shi et al. (Shi et al., 2023) use a genetic algorithm to compress CodeBERT into only 3 MB and reduce its response latency by more than 70%. Overall, the challenge of increasing model sizes and efficient deployment requires further attention from the communities.
@@
Data dependency. In Section 4, we provide a detailed analysis of the datasets used in 229 studies and the data preprocessing process, finding that LLMs rely heavily on a large number of different datasets for training and fine-tuning, posing the data dependency challenge. The quality, diversity, and quantity of data directly affect the performance and generalizability of the models. Given their size, LLMs often require large amounts of data to capture nuances, but obtaining such data can be challenging. Relying on limited or biased datasets may cause the model to inherit these biases, resulting in biased or inaccurate predictions. In addition, the domain-specific data required for fine-tuning can be a bottleneck. Due to the relatively short period of time since the emergence of LLM, such large-scale datasets are still relatively rare, especially in the SE domain. Another issue is the risk of benchmark data contamination, where training and test data overlaps could lead to inflated performance metrics (Zhao et al., 2021). For instance, Brown et al. (Brown et al., 2020) discovered a code bug that prevented them from fully removing all overlapping data. They were unable to afford retraining and resorted to using “cleaned” variants of the benchmarks to mitigate the issue. Moreover, there are grave concerns around the inclusion of Personally Identifiable Information (PII) in pre-training corpora. Instances of PII, such as phone numbers and email addresses, have led to privacy leaks during the prompting process (Kulkarni, 2021; El-Mhamdi et al., 2023).
@@ 8.1.2. Challenges in LLM Generalizability
The generalizability of LLMs refers to the ability of these models to consistently and accurately perform tasks in different tasks, datasets, or domains outside their training environment. While LLMs are trained on massive amounts of data, ensuring extensive knowledge capture, their performance is sometimes problematic when confronted with specific or idiosyncratic tasks outside the scope of their training. This challenge is particularly evident in the SE domain, where we present the application of LLMs to 55 SE tasks in Section 6. We observed that the context and semantics of code or documents vary greatly across projects, languages, or domains. Ensuring that the LLM generalizes well requires careful fine-tuning, validation on different datasets, and continuous feedback loops. Without these measures, models run the risk of over-adapting their training data, thus limiting their usefulness in a variety of real-world applications. Recent studies have shown that the LLMs cannot generalize their good performance to inputs after semantic-preserving transformations. For example, Yang et al. (Yang et al., 2022a) show that the performance of CodeBERT on different tasks decreases significantly after substituting the variables’ names in the input.
@@ 8.1.3. Challenges in LLM Evaluation
We summarized key evaluation metrics used in different types of SE tasks according to four task types: regression, classification, recommendation, and generation (Section 6). We found that when applying LLMs in the software engineering domain, the methodology for evaluating the performance of the models is usually based on a set of predefined metrics. Unfortunately, these metrics (e.g., Accuracy, Recall, or F1-score), while useful in some cases, may not fully capture all the effects and impacts of a model in a given SE task. For example, a model may perform well in terms of accuracy but may fail in processing specific types of inputs or in some specific situations. In addition, these metrics may not capture certain qualitative aspects of the model, such as its interpretability, robustness, or sensitivity to specific types of errors. Some of the most recent studies on LLM4SE tasks (Agrawal et al., 2023; Hu et al., 2023; Singla, 2023; Xu et al., 2023; Yuan et al., 2023a; Zhang et al., 2023e), in which researchers customized some evaluation metrics to assess the performance of models, also further illustrate the limitations of some of the widely used evaluation metrics in the field of LLM.
@@ 8.1.4. Challenges in LLM Interpretability, Trustworthiness, and Ethical Usage
Interpretability and trustworthiness are crucial aspects in the adoption of LLMs for SE tasks. The challenge lies in understanding the decision-making process of these models, as their black-box nature often makes it difficult to explain why or how a particular code snippet or recommendation is generated. Recent studies (Yang et al., 2023b; Wan et al., 2022a; Li et al., 2022c) also show that LLM of code trained on low-quality datasets can have vulnerabilities (e.g., generating insecure code). The lack of interpretability and trustworthiness can lead to uncertainty and hesitation among developers, who may be hesitant to rely on LLM-generated code without a clear understanding of how it was derived. Establishing trust in LLMs requires efforts to develop techniques and tools that provide insights into the model’s internal workings and enable developers to comprehend the reasoning behind the generated outputs. Enhancing interpretability and trustworthiness can ultimately promote the widespread adoption of LLMs in SE, leading to more efficient and effective development practices. Many LLMs are not open and it is unclear what data they have been trained on, both quality and representativeness but also ownership of the source training data. This brings into question ownership of the derivative data, e.g., generated designs, code, or test cases. There is also potential for various adversarial attacks e.g. deliberately seeding LLMs with code vulnerabilities so that automatically generated code snippets have subtle but vulnerable aspects.
@@ 8.2.1. Optimization of LLM4SE
Influence and applications of ChatGPT. ChatGPT’s popularity in recent academic research, as evidenced by its large presence in our 229 analyzed papers, emphasizes its escalating influence and acceptance within academia. Researchers’ preference for ChatGPT over other LLMs and LLM-based applications since its release can be attributed to its computational efficiency, adaptability to various tasks, and potential cost-effectiveness (Laskar et al., 2023; Li et al., 2023b; Xia and Zhang, 2023a). Its applications extend beyond mere code efficiency and debugging, fostering a collaborative era in development. This paradigm shift signifies a broader move towards integrating advanced natural language understanding into conventional coding practices (Laskar et al., 2023; Ma et al., 2023a; Sadik et al., 2023). By thoughtfully analyzing these dynamics and trends, we can foresee the potential pathways for LLMs and LLM applications like ChatGPT in shaping more robust, efficient, and collaborative software development procedures. Such insights stand as a promising indication of the future revolutionary impact of LLMs on SE.
@@
Performance enhancement from task-specific model training. The choice between leveraging commercially available pre-trained models like GPT-4 and building upon open-source frameworks such as LLaMA (Touvron et al., 2023a), Llama 2 (Touvron et al., 2023b), and Alpaca (Alpaca, 2023) (fine-tuned from LLaMA 7B on 52K instruction-following demonstrations) provides a nuanced set of options for individual or organizational customization in specialized tasks. The distinction between these two approaches lies in the degree of control and customization. Pre-trained models like GPT-4 are generally not designed for large-scale retraining due to their proprietary nature, but they allow quick task-specific adaptations with limited data, thereby minimizing computational overhead. On the other hand, frameworks like LLaMA offer an open-source foundation for more extensive customization. While they come pre-trained, organizations often modify the source code and retrain these models on their own large-scale datasets to meet specialized requirements (ymcui, 2023; hiyouga, 2023). This process is computationally intensive, leading to greater resource allocation and cost, but affords the advantage of creating highly domain-specific models. Hence, the primary trade-off is between the ease of use and quick deployment offered by models like GPT-4, and the deep customization capabilities but higher computational demands associated with open-source frameworks like LLaMA.
@@
Collaborative LLMs. From our review it is evident that LLMs have made significant strides in addressing various SE challenges. However, as the complexity of SE tasks continues to grow, there’s an emerging need for more sophisticated and tailored solutions. One promising direction is the concept of Collaborative LLMs. This approach involves integrating multiple LLMs (Dong et al., 2023; Zhao et al., 2023b) or combining LLMs with specialized machine-learning models (Ezzini et al., 2022; Zhang et al., 2022a) to enhance their efficacy for SE tasks. By harnessing the collective strengths of different models, we believe that the SE community can achieve more precise and efficient outcomes, from code completion to bug detection.
@@
Integration of new input forms. In our analysis we observed that the predominant input forms were code-based datasets and text-based datasets. However, there was a noticeable scarcity of graph-based datasets (Kolthoff et al., 2023) (Section 4). Leveraging new input forms of natural language, such as spoken language, diagrams, and multimodal inputs, presents an opportunity to enhance the LLMs’ ability to understand and process diverse user requirements. Integrating spoken language could improve interactions between developers and models, enabling more natural and context-rich communication. Diagrams can facilitate visual representations of code and requirements, offering a complementary perspective for code generation. Furthermore, multimodal inputs that combine text, audio, and visual cues could offer a more comprehensive context understanding, leading to more accurate and contextually appropriate code generation. Additionally, exploring graph-based datasets could be crucial for addressing complex code scenarios, as graphs capture the structural relationships and dependencies in code, allowing LLMs to better comprehend code interactions and dependencies.
Widening LLM applications across SE phases. We observed a pronounced emphasis on the application of LLMs in software development and maintenance. These areas have undoubtedly benefited from the capabilities of LLMs, leading to enhanced code completion (Izadi et al., 2022; Li et al., 2022e; Liu et al., 2023e), bug detection (Ciborowska and Damevski, 2023; Feng and Chen, 2023; Kang et al., 2023), and other related tasks. The current application of LLMs in requirements engineering, software design, and software management remains relatively sparse. This presents a significant opportunity: by expanding the use of LLMs to these under-explored areas, we can potentially improve how requirements are elicited, how software designs are conceptualized, and how projects are managed.

### v6

@@ 6.7. How are LLMs used in software management?
RQ4 - Summary
(1) We categorized SE tasks into six activities: requirements engineering, software design, software development, software quality assurance, software maintenance, and software management. Subsequently, we summarized the specific applications of LLMs in these SE activities. (2) We identified a total of 85 SE tasks and found that LLMs are most widely used in software development, with 229 papers mentioning over 24 SE tasks. The least applied area, software management, was mentioned in only three studies. (3) Code generation and program repair are the most prevalent tasks for employing LLMs in software development and maintenance activities. We analyze the top-performing LLMs repeatedly validated in these tasks and summarize novel findings.
@@ 7. Threats to Validity
Paper search omission. One key limitation is the possibility of omitting relevant papers during the search process. When gathering papers related to LLM4SE tasks from various publishers, it is possible to miss some papers due to incomplete summarization of keywords for software engineering tasks or LLMs. To address this concern, we adopted a comprehensive approach, combining manual search, automated search, and snowballing techniques, to minimize the risk of missing relevant papers. For manual search, we systematically searched for LLM papers related to SE tasks in six top-tier SE venues and extracted authoritative and comprehensive SE tasks and LLM keywords from these sources. Using these constructed search strings, we conducted automated searches on seven widely used publisher platforms. Additionally, to further augment our search results, we employed both forward and backward snowballing.
Study selection bias. Another limitation is the potential study selection bias. We established inclusion and exclusion criteria to perform the initial selection of papers, followed by manual verification based on quality assessment criteria (QAC). This process involves a combination of automated and manual procedures. The automated selection process may result in mislabeling of papers due to incomplete or ambiguous information in their corresponding BibTeX records. To mitigate this issue, any papers that cannot be confidently excluded are temporarily retained for manual verification. However, the manual verification stage could be influenced by the subjective judgment and biases of the researchers, affecting the accuracy of the quality assessment of papers. To address these concerns, we invited two experienced reviewers in the fields of SE and LLM research to conduct a secondary review of the study selection results. This step aims to enhance the accuracy of our paper selection and minimize the likelihood of omission or misclassification. By using these measures, we strive to ensure that the selected papers are accurate and comprehensive, minimizing the impact of study selection bias and enhancing the reliability of our systematic literature review. We additionally provide a replication package [^6] https://github.com/xinyi-hou/LLM4SE_SLR for others to view.
Empirical knowledge bias. This SLR, along with 395 relevant studies in the LLM4SE field, answers four RQs. This implies the need for manual analysis and understanding of each study. In this process, there may be biases introduced by subjective judgments and experiential knowledge. To minimize potential errors in this regard, we have made the following efforts. Firstly, in determining the RQs, as the first comprehensive overview of the LLM4SE field, we aim to provide a comprehensive interpretation of the current state and trends in this domain. Considering the commonality in AI4SE research, we referred to Yang et al.’s survey on DL4SE (Yang et al., 2022b) during our RQ formulation. We finally decided to focus on LLM types, datasets, tuning, evaluation, and targeted SE tasks. Secondly, for the understanding and analysis of each study, to ensure accurate comprehension of paper details, before addressing each RQ, we extensively reviewed relevant literature to predefine the approximate categories and details for each RQ. For example, in RQ3, based on prior work (Sahoo et al., 2024; Weyssow et al., 2023a; Zhao et al., 2023d), we identified differences between tuning techniques for LLMs and those commonly used in traditional machine learning, such as prompt engineering and PEFT.
@@ 8.1. Challenges
Model size and deployment. The size of LLMs has seen a marked increase over time, moving from GPT-1’s 117M parameters to GPT-2’s 1.5B, and further to GPT-3’s 175B parameters (Yang et al., 2023b). The billions and even trillions (Moss, 2021) of parameters pose significant storage, memory, and computational challenges, which can hinder LLMs in resource-limited and real-time scenarios, especially when developers lack access to powerful GPUs or TPUs. CodeBERT (Feng et al., 2020), a pre-trained model proposed in 2019, has a total of 125M parameters, resulting in a large model size of 476 MB. Recently proposed models like Codex (Chen et al., 2021b) and CodeGen (Nijkamp et al., 2022a), have over 100 billion parameters and over 100 GB in size. The large sizes also require more computational resources. As pointed out by Hugging Face team (Bekman, 2022), training a 176B model (i.e., BLOOM (Scao et al., 2022)) on 1.5 TB datasets consumes an estimated 1,082,880 GPU hours. Similarly, the training of the GPT-NeoX-20B model (Black et al., 2022) on the Pile dataset (Gao et al., 2020), encompassing over 825 GiB of raw text data, requires the deployment of eight NVIDIA A100-SXM4-40GB GPUs. Each of these GPUs comes with a price tag of over 6,000 dollars (Amazon, 2023b), and the training extends to 1,830 hours or approximately 76 days. Moreover, even training a relatively smaller model like the PolyCoder (2.7B) (Xu et al., 2022), employing eight NVIDIA RTX 8000 GPUs on a single machine, demands a commitment of around 6 weeks. These examples illustrate the significant computational costs associated with training LLMs. These also have significant energy costs with predictions of massively increased energy usage by LLM-based platforms (Rillig et al., 2023). Fortunately, there are preliminary studies on reducing code models’ size and improving their efficiency. Shi et al. (Shi et al., 2023b) use a genetic algorithm to compress CodeBERT into only 3 MB and reduce its response latency by more than 70%. Overall, the challenge of increasing model sizes and efficient deployment requires further attention from the communities.
@@
Data dependency. In Section 4, we provide a detailed analysis of the datasets used in 395 studies and the data preprocessing process, finding that LLMs rely heavily on a large number of different datasets for training and fine-tuning, posing the data dependency challenge. The quality, diversity, and quantity of data directly affect the performance and generalizability of the models. Given their size, LLMs often require large amounts of data to capture nuances, but obtaining such data can be challenging. Relying on limited or biased datasets may cause the model to inherit these biases, resulting in biased or inaccurate predictions. In addition, the domain-specific data required for fine-tuning can be a bottleneck. Due to the relatively short period of time since the emergence of LLM, such large-scale datasets are still relatively rare, especially in the SE domain. Another issue is the risk of benchmark data contamination, where training and test data overlaps could lead to inflated performance metrics (Zhao et al., 2021). For instance, Brown et al. (Brown et al., 2020) discovered a code bug that prevented them from fully removing all overlapping data. They were unable to afford retraining and resorted to using “cleaned” variants of the benchmarks to mitigate the issue. Moreover, there are grave concerns around the inclusion of Personally Identifiable Information (PII) in pre-training corpora. Instances of PII, such as phone numbers and email addresses, have led to privacy leaks during the prompting process (Kulkarni, 2021; El-Mhamdi et al., 2023).
@@ 8.1.2. Challenges in LLM Generalizability
The generalizability of LLMs refers to the ability of these models to consistently and accurately perform tasks in different tasks, datasets, or domains outside their training environment. While LLMs are trained on massive amounts of data, ensuring extensive knowledge capture, their performance is sometimes problematic when confronted with specific or idiosyncratic tasks outside the scope of their training. This challenge is particularly evident in the SE domain, where we present the application of LLMs to 85 SE tasks in Section 6. We observed that the context and semantics of code or documents vary greatly across projects, languages, or domains. Ensuring that the LLM generalizes well requires careful fine-tuning, validation on different datasets, and continuous feedback loops. Without these measures, models run the risk of over-adapting their training data, thus limiting their usefulness in a variety of real-world applications. Recent studies have shown that the LLMs cannot generalize their good performance to inputs after semantic-preserving transformations. For example, Yang et al. (Yang et al., 2022a) show that the performance of CodeBERT on different tasks decreases significantly after substituting the variables’ names in the input.
@@ 8.1.3. Challenges in LLM Evaluation
We summarized key evaluation metrics used in different types of SE tasks according to four task types: regression, classification, recommendation, and generation (Section 6). We found that when applying LLMs in the software engineering domain, the methodology for evaluating the performance of the models is usually based on a set of predefined metrics. Unfortunately, these metrics (e.g., Accuracy, Recall, or F1-score), while useful in some cases, may not fully capture all the effects and impacts of a model in a given SE task. For example, a model may perform well in terms of accuracy but may fail in processing specific types of inputs or in some specific situations. In addition, these metrics may not capture certain qualitative aspects of the model, such as its interpretability, robustness, or sensitivity to specific types of errors. Some of the most recent studies on LLM4SE tasks (Agrawal et al., 2023; Hu et al., 2023; Singla, 2023; Xu et al., 2023b; Yuan et al., 2023a; Zhang et al., 2023l), in which researchers customized some evaluation metrics to assess the performance of models, also further illustrate the limitations of some of the widely used evaluation metrics in the field of LLM.
@@ 8.1.4. Challenges in LLM Interpretability, Trustworthiness, and Ethical Usage
Interpretability and trustworthiness are crucial aspects in the adoption of LLMs for SE tasks. The challenge lies in understanding the decision-making process of these models, as their black-box nature often makes it difficult to explain why or how a particular code snippet or recommendation is generated. Recent studies (Yang et al., 2023e; Wan et al., 2022a; Li et al., 2022c) also show that LLM of code trained on low-quality datasets can have vulnerabilities (e.g., generating insecure code). The lack of interpretability and trustworthiness can lead to uncertainty and hesitation among developers, who may be hesitant to rely on LLM-generated code without a clear understanding of how it was derived. Establishing trust in LLMs requires efforts to develop techniques and tools that provide insights into the model’s internal workings and enable developers to comprehend the reasoning behind the generated outputs. Enhancing interpretability and trustworthiness can ultimately promote the widespread adoption of LLMs in SE, leading to more efficient and effective development practices. Many LLMs are not open and it is unclear what data they have been trained on, both quality and representativeness but also ownership of the source training data. This brings into question ownership of the derivative data, e.g., generated designs, code, or test cases. There is also potential for various adversarial attacks e.g. deliberately seeding LLMs with code vulnerabilities so that automatically generated code snippets have subtle but vulnerable aspects.
@@ 8.2.1. Optimization of LLM4SE
Influence and applications of ChatGPT. ChatGPT’s popularity in recent academic research, as evidenced by its large presence in our 395 analyzed papers, emphasizes its escalating influence and acceptance within academia. Researchers’ preference for ChatGPT over other LLMs and LLM-based applications since its release can be attributed to its computational efficiency, adaptability to various tasks, and potential cost-effectiveness (Laskar et al., 2023; Li et al., 2023c; Xia and Zhang, 2023a). Its applications extend beyond mere code efficiency and debugging, fostering a collaborative era in development. This paradigm shift signifies a broader move towards integrating advanced natural language understanding into conventional coding practices (Laskar et al., 2023; Ma et al., 2023a; Sadik et al., 2023). By thoughtfully analyzing these dynamics and trends, we can foresee the potential pathways for LLMs and LLM applications like ChatGPT in shaping more robust, efficient, and collaborative software development procedures. Such insights stand as a promising indication of the future revolutionary impact of LLMs on SE.
@@
Performance enhancement from task-specific model training. The choice between leveraging commercially available pre-trained models like GPT-4 and building upon open-source frameworks such as Llama 2 (Touvron et al., 2023b), Gemma (Google, 2024), and Mistral (AI, 2023) provides a nuanced set of options for individual or organizational customization in specialized tasks. The distinction between these two approaches lies in the degree of control and customization. Pre-trained models like GPT-4 are generally not designed for large-scale retraining due to their proprietary nature, but they allow quick task-specific adaptations with limited data, thereby minimizing computational overhead. On the other hand, frameworks like LLaMA offer an open-source foundation for more extensive customization. While they come pre-trained, organizations often modify the source code and retrain these models on their own large-scale datasets to meet specialized requirements (ymcui, 2023; hiyouga, 2023). This process is computationally intensive, leading to greater resource allocation and cost, but affords the advantage of creating highly domain-specific models. Hence, the primary trade-off is between the ease of use and quick deployment offered by models like GPT-4, and the deep customization capabilities but higher computational demands associated with open-source frameworks like LLaMA.
@@
Collaborative LLMs. From our review it is evident that LLMs have made significant strides in addressing various SE challenges. However, as the complexity of SE tasks continues to grow, there’s an emerging need for more sophisticated and tailored solutions. One promising direction is the concept of Collaborative LLMs. This approach involves integrating multiple LLMs (Dong et al., 2023b; Zhao et al., 2023b) or combining LLMs with specialized machine-learning models (Ezzini et al., 2022; Zhang et al., 2022a) to enhance their efficacy for SE tasks. By harnessing the collective strengths of different models, we believe that the SE community can achieve more precise and efficient outcomes, from code completion to bug detection.
@@
Integration of new input forms. In our analysis, we observed that the predominant input forms were code-based datasets and text-based datasets. However, there was a noticeable scarcity of graph-based datasets (Kolthoff et al., 2023) (Section 4). Leveraging new input forms of natural language, such as spoken language, diagrams, and multimodal inputs, presents an opportunity to enhance the LLMs’ ability to understand and process diverse user requirements. Integrating spoken language could improve interactions between developers and models, enabling more natural and context-rich communication. Diagrams can facilitate visual representations of code and requirements, offering a complementary perspective for code generation. Furthermore, multimodal inputs that combine text, audio, and visual cues could offer a more comprehensive context understanding, leading to more accurate and contextually appropriate code generation. Additionally, exploring graph-based datasets could be crucial for addressing complex code scenarios, as graphs capture the structural relationships and dependencies in code, allowing LLMs to better comprehend code interactions and dependencies.
Widening LLM applications across SE phases. We observed a pronounced emphasis on the application of LLMs in software development and maintenance. These areas have undoubtedly benefited from the capabilities of LLMs, leading to enhanced code completion (Izadi et al., 2022; Li et al., 2022e; Liu et al., 2023l), bug detection (Ciborowska and Damevski, 2023; Feng and Chen, 2023; Kang et al., 2023b), and other related tasks. The current application of LLMs in requirements engineering, software design, and software management remains relatively sparse. This presents a significant opportunity: by expanding the use of LLMs to these under-explored areas, we can potentially improve how requirements are elicited, how software designs are conceptualized, and how projects are managed.